### v1

@@ 5.1 Filter Augmented Input
To investigate and assess the impact of the proposed Filter Augmented Input, we conduct experiments on each traditional feature descriptor discussed in the Related Work, within the framework of our proposed MSFA method. The findings, detailed in Table 3(a), indicate that incorporating these handcrafted features notably enhances the performance of the detector. Additionally, our analysis reveals that converting image pixels into handcrafted feature spaces significantly minimizes the distributional gaps between the ImageNet and the SARDet-100K datasets. This is particularly evident in the Pearson Correlation Coefficients (PCC) between the inputs of the ImageNet and SARDet-100K datasets, as illustrated in Table 3(b). This underscores the efficacy of the proposed approach in bridging the domain gap between natural and SAR images, thereby enhancing the efficiency of knowledge transfer from the pretraining process.
@@
Table 3: (a): Comparison of different Filter Augmented Inputs using Faster R-CNN and Res50 as the detection model. (b): Pearson Correlation Coefficients (PCC) of ImageNet and SARDet-100k on RGB and handcraft feature spaces. PCC increase dramatically from pixel space to WST feature space, as illustrated in Fig. 3(a).
[…]
  SAR (as RGB) 50.2 83.0
SAR+Canny 50.7 83.6
SAR+Hog 50.7 83.5
SAR+Haar 50.6 83.4
SAR+WST 51.1 83.9
SAR+GRE 50.6 83.8
SAR+Hog+Haar+WST 51.1 84.0
Domain   PCC $\uparrow$
  Pixel Space   0.394
Canny Space   0.992
Hog Space   0.995
Haar Space   0.990
WST Space   0.996
GRE Space   0.984
@@
Table 4: Comparison of different pretrain strategies using Faster-RCNN and Res50 as the detection model.
@@
5.2 Multi-stage Pretrain
To evaluate the effectiveness of the proposed multi-stage pretraining approach, we conduct experiments in which we keep the input modality consistent and finetuned the detection model using various pretraining strategies on the SARDet-100K dataset. As a baseline, Exp. 1 takes single-channel SAR data as input, pretrains the backbone model on ImageNet for 100 epochs, and then directly finetune the detector on the SARDet-100K dataset (following the widely used default setting). In addition to the baseline, we perform a second stage pretrain specifically for object detection on optical remote sensing datasets, such as DOTA [73] or DIOR [28]. (Details on DOTA and DIOR datasets can be found in the Supplementary Materials). Following the second pretraining stage, we finetune the model either solely on the backbone or on the entire framework.
In summary, our proposed multi-stage pretraining strategy in MSFA alleviates both the data domain gap and the model gap between pretraining and the downstream model, leading to significant enhancements in SAR detection performance. Detailed experimental results and visualization are available in the Supplementary Materials.
Figure 6: Generalization of MSFA on different detection frameworks (a) and different backbones (b). INP: Traditional ImageNet Pretrain on backbone network only.
@@ 5.3 Generalizability of MSFA
To assess the effectiveness and generalizability of the proposed MSFA, we conduct experiments using various detectors and backbones, as presented in Fig. 5(a) and 5(b). Significant performance improvement is observed across different frameworks (including single-stage [53, 4, 44], two-stage [34, 31, 59], and end2end [93, 58, 40]) and diverse backbones (including ResNets [24], ConvNexts [43], VANs [23], and Swin-Transformer [42] networks). It provides strong evidence for the effectiveness and wide applicability of our proposed method. Furthermore, we observe stable performance improvements as we scale up the backbone size, as shown in Fig. 5(b), indicating the good scalability of our proposed method.
@@
5.4 Comparasion with SOTAs
We compare various SOTA methods, including both general object detection models [44, 53, 14, 64, 4, 86, 72], as well as SAR object detection models [7, 26, 88, 13, 70, 41, 38, 74, 22, 21]. We evaluate their performance on the SSDD and HRSID datasets, which are the commonly used benchmarks for SAR object detection. To leverage the superior efficiency and performance of the VAN [23] backbone (as shown in Fig. 5(b)), we employ the classic faster R-CNN detection framework with the lightweight VAN-B (Param. 26.6M) backbone as our detection model. The results presented in Table 5 demonstrate that our MSFA method outperforms all the compared methods by a significant margin. Specifically, MSFA achieves a mAP@50 of 97.9% on the SSDD dataset and a mAP@50 of 83.7% on the HRSID dataset, setting new state-of-the-art results. It is noteworthy that our method is the only open-sourced method among the compared SAR detection SOTAs.
@@ 5.4 Comparasion with SOTAs
  General Detectors Grid R-CNN [44] 88.9 79.4
Faster R-CNN [53] 89.7 80.7
Cascade R-CNN [4] 90.5 81.3
Free-Anchor [86] 91.0 81.8
Double-Head R-CNN [72] 91.1 82.1
PANET [41] 91.2 81.6
DCN [14] 92.3 82.1
SAR Detectors NNAM [7] 79.8 -
DCMSNM [26] 89.6 -
ARPN [88] 89.9 81.8
DAPN [13] 90.6 81.8
HR-SDNet [70] 90.8 81.5
SER Faster R-CNN [38] 91.5 81.5
FBR-Net [21] 94.1 -
CenterNet++ [22] 95.1 -
CRTransSar [74] 97.0 -
@@ 7 Conclusion
Acknowledgement
We extend our deepest gratitude to Bo Zhang, Chenglong Li, Tian Tian, Tianwen Zhang, Xiaoling Zhang (ordered alphabetically by first name) and numerous other researchers for permitting us to integrate their datasets. Their contributions have significantly advanced and promoted research in this field.
This work is supported by the Young Scientists Fund of the National Natural Science Foundation of China (No.62206134), National Science Fund of China (No. 62361166670 and No. 62276145), the Fundamental Research Funds for the Central Universities 070-63233084, and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). Computation is supported by the Supercomputing Center of Nankai University (NKSC).
8 Appendix
8.1 SARDet-100K Dataset Visualization
@@
8.2 Multi-Stage Self-Multimodal
Fig. S8 provides a visual illustration of the proposed self-multimodal data input. It involves concatenating the original single-channel grayscale Synthetic Aperture Radar (SAR) image, denoted as $x$, with the self-multimodal representation $M_{i}^{x}$.
Figure S8: Illustration of Self-Multimodal input.
8.3 Experimental Results
Table S6 serves as an extension of Table 3 from the main paper, showcasing a comparison of various self-multimodal inputs utilizing Faster R-CNN with Resnet-50 as the detection model.
@@ 8.3 Experimental Results
8.4 Implementaion Details
For the second stage of the proposed multistage pretrain, we select the large-scale optical remote sensing DOTA dataset as the main dataset. We also do comparison experiments on the DIOR dataset to find out the downstream performance affection factors on the second stage pretrain dataset. Their detailed information is illustrated in Table S9. For the DOTA dataset, we perform an extra dataset preprocess because DOTA dataset images have a large range of different image resolutions. The original image contains only 1,411 training images. To have more image and multiscale instances for effective training, we follow [32] to adopt a multi-scale dataset splitting strategy by rescaling the original high-resolution images into three distinct scales (x0.5, x1.0, x1.5), and then cropping each scaled image into 1024×1024 patches with an overlap of each patch of 500 pixels, to avoid destruction division on the instance at patch borders. With the preprocessed dataset, load resize as 1024*1024, with a RandomFlip probability of 0.5. For the DIOR dataset, we train the model by resizing the images as 800*800, with a RandomFlip probability of 0.5. For fine-tuning on SARDet-100k, SSDD and HRSID, we train by resizing the image into 800*800, with a RandomFlip probability of 0.5.
We primarily conduct our experiments using the MMPretrain [11] and the MMDetection [8] frameworks. For detailed information on the hyperparameters and training settings, please refer to Table S10.
@@ 8.4 Implementaion Details
Table S10: Hyper-parameter of pretrain and finetune settings.
@@
8.5 Detection Result Visualizations
@@
[Missing detection]Refer to caption [False detection]Refer to caption [Inaccurate localization]Refer to caption

### v2

@@ 5.1 Filter Augmented Input
To investigate and assess the impact of the proposed Filter Augmented Input, we conduct experiments on each traditional feature descriptor discussed in the Related Work, within the Input mAP $\uparrow$ mAP50 $\uparrow$   SAR (as RGB) 50.2 83.0 SAR+Canny 50.7 83.6 SAR+Hog 50.7 83.5 SAR+Haar 50.6 83.4 SAR+WST 51.1 83.9 SAR+GRE 50.6 83.8 SAR+Hog+Haar+WST 51.1 84.0 Table 3: Comparison of different Filter Augmented Inputs using Faster R-CNN and ResNet50 as the detection model. Domain   PCC $\uparrow$   Pixel Space   0.394 Canny Space   0.992 Hog Space   0.995 Haar Space   0.990 WST Space   0.996 GRE Space   0.984 Table 4: Pearson Correlation Coefficients (PCC) of ImageNet and SARDet-100k on RGB and handcrafted feature spaces.
framework of our proposed MSFA method. The findings, detailed in Table 5.1, indicate that incorporating these handcrafted features notably enhances the performance of the detector. Additionally, our analysis reveals that converting image pixels into handcrafted feature spaces significantly minimizes the distributional gaps between the ImageNet and the SARDet-100K datasets. This is particularly evident in the Pearson Correlation Coefficients (PCC) between the inputs of the ImageNet and SARDet-100K datasets, as illustrated in Table 5.1. This underscores the efficacy of the proposed approach in bridging the domain gap between natural and SAR images, thereby enhancing the efficiency of knowledge transfer from the pretraining process.
@@
5.2 Multi-stage Pretrain
To evaluate the effectiveness of the proposed multi-stage pretraining approach, we conduct experiments in which we keep the input modality consistent and finetuned the detection model using various pretraining strategies on the SARDet-100K dataset. As a baseline, Exp. 1 takes single-channel SAR data as input, pretrains the backbone model on ImageNet for 100 epochs, and then directly finetune the detector on the SARDet-100K dataset (following the widely used default setting). In addition to the baseline, we perform a second stage pretrain specifically for object detection on optical remote sensing datasets, such as DOTA [73] or DIOR [27]. (Details on DOTA and DIOR datasets can be found in the Appendix). Following the second pretraining stage, we finetune the model either solely on the backbone or on the entire framework.
In summary, our proposed multi-stage pretraining strategy in MSFA alleviates both the data domain gap and the model gap between pretraining and the downstream model, leading to significant enhancements in SAR detection performance. Detailed experimental results and visualization are available in the Appendix.
Table 5: Comparison of different pretrain strategies using Faster-RCNN and ResNet50 as the detection model.
@@ 5.2 Multi-stage Pretrain
Figure 4: Generalization of MSFA on different detection frameworks (a) and different backbones (b). Models are finetuned and tested on SARDet-100K dataset. INP: Traditional ImageNet Pretrain on backbone network only.
@@ 5.3 Generalizability of MSFA
To assess the effectiveness and generalizability of the proposed MSFA, we conduct experiments using various detectors and backbones, as presented in Fig. 4(a) and 4(b). Significant performance improvement is observed across different frameworks (including single-stage [53; 4; 43], two-stage [33; 30; 59], and end2end [94; 58; 39]) and diverse backbones (including ResNets [23], ConvNexts [42], VANs [22], and Swin-Transformer [41] networks). It provides strong evidence for the effectiveness and wide applicability of our proposed method. Furthermore, we observe stable performance improvements as we scale up the backbone size, as shown in Fig. 4(b), indicating the good scalability of our proposed method.
@@
5.4 Comparison with SOTAs
We compare various SOTA methods, including both general object detection models [43; 53; 14; 64; 4; 87; 72], as well as SAR object detection models [7; 25; 89; 13; 70; 40; 37; 74; 21; 20]. We evaluate their performance on the SSDD and HRSID datasets, which are the commonly used benchmarks for SAR object detection. To leverage the superior efficiency and performance of the VAN [22] backbone (as shown in Fig. 4(b)), we employ the classic faster R-CNN detection framework with the lightweight VAN-B (Param. 26.6M) backbone as our detection model. The results presented in Table 6 demonstrate that our MSFA method outperforms all the compared methods by a significant margin. Specifically, MSFA achieves a mAP@50 of 97.9% on the SSDD dataset and a mAP@50 of 83.7% on the HRSID dataset, setting new state-of-the-art results. It is noteworthy that our method is the only open-sourced method among the compared SAR detection SOTAs.
@@ 5.4 Comparison with SOTAs
  General Detectors Grid R-CNN [43] 2019 88.9 79.4
Faster R-CNN [53] 2015 89.7 80.7
Cascade R-CNN [4] 2019 90.5 81.3
Free-Anchor [87] 2019 91.0 81.8
Double-Head R-CNN [72] 2020 91.1 82.1
PANET [40] 2018 91.2 81.6
DCN [14] 2017 92.3 82.1
SAR Detectors NNAM [7] 2019 79.8 -
DCMSNM [25] 2018 89.6 -
ARPN [89] 2020 89.9 81.8
DAPN [13] 2019 90.6 81.8
HR-SDNet [70] 2020 90.8 82.5
SER Faster R-CNN [37] 2018 91.5 81.5
FBR-Net [20] 2020 94.1 -
NRENet [44] 2024 94.6 75.6
CenterNet++ [21] 2021 95.1 -
CRTransSar [74] 2022 97.0 -
SARATR-X [77] 2024 97.3 80.3
@@ 7 Conclusion
Our research endeavours to overcome the current obstacles prevalent in SAR object detection. We anticipate our contributions will pave the way for future research and innovations in this domain.
8 Acknowledgement
We extend our deepest gratitude to the following researchers, listed alphabetically by first name: Hong Zhang, Runfan Xia, Shunjun Wei, Tianwen Zhang, Xian Sun, Xiaofang Zhu, Xiaoling Zhang, and other contributing researchers, for allowing us to integrate their datasets. Their contributions have greatly advanced and promoted research in this field.
This work is supported by the National Science Fund of China (No. 62361166670 and No. 62276145), the Young Scientists Fund of the National Natural Science Foundation of China (No.62206134), the Fundamental Research Funds for the Central Universities 070-63233084, and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). Computation is supported by the Supercomputing Center of Nankai University (NKSC).
Appendix A Appendix
A.1 SARDet-100K Dataset Visualization
@@
Figure S6: (a) SARDet-100K dataset standardization process, encompassing set splitting, large image slicing, and label annotation format unification. (b) Percentage of instances for each category and average instance area (in pixels) in SARDet-100K.
A.2 Standardization.
We collect a total of 10 publicly available high-quality datasets that are not only diverse but also have no conflicting object categories. These data are released by or collected from different countries and institutions, such as scientific research departments in China, space departments in Europe, and military departments in the United States. To ensure consistency across the collected datasets, it is necessary to undertake standardization processes. This involves addressing variations in train-val-test splitting status, image resolutions, and annotation formats. The overview of SARDet-100K dataset process pipeline is illustrated in Fig. S6(a).
Furthermore, we convert all dataset annotations into the COCO annotation format [34]. This step ensures consistency and compatibility among the different datasets. Consequently, the merged dataset, SARDet-100K, is also standardized in the COCO format, which is readily compatible with popular open-source detection code frameworks, eliminating the need for additional manual data preprocessing. Fig. S6(b) provides an overview of the category-level statistics for the SARDet-100K dataset.
A.3 Handcrafted Feature Descriptors
Histogram of Oriented Gradients [15] (HOG). HOG is a widely used local feature descriptor in image processing and computer vision. It captures and represents the image’s local structure and shape information by analyzing the distribution of gradient orientations. HOG is proven to be effective for SAR image classification and object detection tasks because it is invariant to random noise in SAR images. Early works employ HOG for SAR object recognition [56; 49], and recent studies demonstrate the effectiveness of HOG features for SAR image classification [83; 85], as well as the model pretraining [66] in modern neural networks.
Canny Edge Detector [5]. Canny is a widely used edge detection algorithm, which aims to identify significant edges in an image while minimizing noise and spurious responses. The algorithm utilizes Gaussian smoothing, pixel orientation gradient magnitude, and non-maximum suppression. Early research [28; 38] recognize the advantage of Canny for SAR image processing. Recent works also verify the effectiveness of the Canny feature in SAR image edge detection [69], interference detection [18], and image registration [86]. Other than Canny Edge Detector, Gradient by Ratio Edge (GRE) [29], a recently proposed edge detector, leverages SAR-HOG [16] and SAR-SIFT [56] to achieve effective edge detection in SAR images.
Haar-like [62] Feature Descriptor. Haar-like features are commonly used for face detection [46; 2], pedestrian detection [81], and other target detection tasks [32; 36]. They describe the characteristics of an image by utilizing predefined feature templates. Haar-like features can capture linear features, edge features, point features, and diagonal features. Early studies [54; 55] demonstrate the robustness of Haar-like features for SAR object detection. A recent approach called MSRIHL [1] integrates low-level Haar-like features into deep learning models for accurate SAR object detection, highlighting its potential effectiveness.
Wavelet Scattering Transform [45] (WST). The WST is a powerful signal-processing technique that is widely used in image processing. It aims to extract robust and discriminative low-level and high-level features simultaneously. By capturing both high-frequency and low-frequency information, it provides a rich representation of both local and global image features. The hierarchical representation offered by the Wavelet Scattering Transform enables feature extraction at different scales and resolutions, which is particularly useful for capturing fine details of small objects in SAR images and robustly handling high-frequency noise. Vidal et al. [61] conduct a comprehensive investigation demonstrating the high potential performance of Wavelet Transform in SAR image denoising. Many recent works [93; 26; 50; 78] utilize Wavelet Transform or WST for robust feature extraction and integration with CNN networks for target recognition. These studies validate the feasibility of incorporating WST features into modern CNN models.
Figure S7: Visualization of handcrafted features on SAR images. (To facilitate visualization, the features are average pooled and represented as a single channel.
A.4 Multi-Stage Filter-Augmentation
Fig. S8 provides a visual illustration of the proposed filter-augmented data input. It involves concatenating the original single-channel grayscale Synthetic Aperture Radar (SAR) image, denoted as $x$, with the filter-augmentation representation $M_{i}^{x}$.
Figure S8: Illustration of filter-augmented input.
@@ A.4 Multi-Stage Filter-Augmentation
A.5 Experimental Results
A.5.1 Main results on SARDet-100K.
Table S7 serves as an extension of Table 3 from the main paper, showcasing a comparison of various filter-augmented inputs utilizing Faster R-CNN with ResNet-50 as the detection model.
@@ A.5 Experimental Results
To ensure a fair comparison, we evaluate MSFA and vanilla detection models under similar computational budgets. In Table S11, we train Faster-RCNN models (with a ResNet-50 ImageNet-1K pretrained backbone) with 12 epochs of MSFA pretraining on the DOTA dataset, followed by finetuning on the SARDet-100K dataset. This is compared to directly finetuning the model for 36 epochs on the SARDet-100K dataset. With the same total number of training epochs and a slightly less total number of iterations, the proposed MSFA achieves a significantly higher (1.7% higher) mAP result on the SARDet-100K test set. MSFA pretraining incorporates general knowledge from both natural and remote sensing datasets, which allows efficient and effective finetuning and helps mitigate overfitting in downstream tasks. It is also important to note that, MSFA pretraining is a one-time effort. The pretrained MSFA models can be reused for finetuning different SAR detection datasets.
Table S11: Comparison of MSFA and Finetuning Performance under similar computational budgets. INP: Backbones are pretrained on ImageNet.
Model INP MSFA Epoch Finetune Epochs Total Epochs Total Iterations mAP
  Faster-RCNN [53] 12 24 36 16.1k 54.5
Faster-RCNN [53] 0 36 36 17.7k 52.8
A.5.2 Comparison of SARDet-100K with other datasets.
To assess the quality of the proposed SARDet-100K dataset as a large-scale SAR object detection benchmark, we evaluate different models on this dataset and compare the results with those obtained from other popular benchmark datasets, such as SSDD [84] and HRSID [71]. The results are presented in Fig S9. These results indicate that the performance of modern models on our dataset has not yet reached saturation. Among the various models reviewed, there is an 8.4% performance gap between the weakest and strongest models, whereas, for SSDD and HRSID, this gap is only 4.1% and 4.3%, respectively. This suggests that SSDD and HRSID are relatively simple for most existing models, leading to performance saturation on these smaller datasets.
Furthermore, we observed that on these smaller datasets, models with a relatively large number of parameters tend to suffer from performance degradation due to overfitting. For example, ResNet-152 underperforms compared to ResNet-101 on SSDD, and ResNet-101 underperforms compared to ResNet-50 and ResNet-18 on HRSID. However, on the large-scale SARDet-100K, this issue does not arise. On our larger dataset, models continue to benefit from increased model size, indicating that our proposed dataset is suitable for developing relatively larger models for large-scale SAR object detection.
Figure S9: Evaluate different backbone models on SARDet-100K dataset and other previous popular benchmarks (SSDD [84] and HRSID [71]). Backbones are plugged under Faster-RCNN [53] detection framework.
A.6 Implementaion Details
For the second stage of the proposed multistage pretrain, we select the large-scale optical remote sensing DOTA dataset as the main dataset. We also do comparison experiments on the DIOR dataset to find out the downstream performance affection factors on the second stage pretrain dataset. Their detailed information is illustrated in Table S10.
For the DOTA dataset, we perform an extra dataset preprocess because DOTA dataset images have a large range of different image resolutions. The original image contains only 1,411 training images. To have more image and multiscale instances for effective training, we follow [31] to adopt a multi-scale dataset splitting strategy by rescaling the original high-resolution images into three distinct scales (x0.5, x1.0, x1.5), and then cropping each scaled image into 1024×1024 patches with an overlap of each patch of 500 pixels, to avoid destruction division on the instance at patch borders. With the preprocessed dataset, load resize as 1024*1024, with a RandomFlip probability of 0.5. For the DIOR dataset, we train the model by resizing the images as 800*800, with a RandomFlip probability of 0.5.
For finetuning on SARDet-100k, SSDD and HRSID, we train by resizing the image into 800*800, with a RandomFlip probability of 0.5. We train the model by 12 epochs on the train set and test the model on the test set with the 12th epoch checkpoint.
We primarily conduct our experiments using the MMPretrain [11] and the MMDetection [8] frameworks, on 8 RTX-3090 GPUs (24G). For detailed information on the hyperparameters and training settings, please refer to Table S12.
Table S12: Hyper-parameter of pretrain and finetune settings. Cls.: Classification, Det.: Detection, B.S.: Batch Size, L.R.: Learning Rate.
@@ A.6 Implementaion Details
A.7 Detection Result Visualizations
@@ A.8 Failure Scenarios
Regarding fine-grained category detection performance, Table S13 presents the detection results of Faster-RCNN [53] with ConvNext-B [42] and MSFA pretraining. Notably, the model exhibits relatively low performance for objects with:
  • Small sizes, such as Tank (with an average area of 461 pixels)
  • Large length-width ratios, such as Bridge
  • High appearance variability, such as Airplane
However, the primary focus of this work is to address the domain and model gaps existing in SAR object detection pretraining and finetuning. Our method demonstrates excellent compatibility with most existing deep networks and can seamlessly integrate with models specifically designed to tackle the above challenging scenarios.
@@
Table S13: ConvNext-B MSFA
Category mAP @50 @75 @s @m @l
  ship 0.669 0.923 0.783 0.653 0.714 0.555
aircraft 0.455 0.753 0.466 0.419 0.458 0.47
car 0.655 0.985 0.792 0.553 0.674 n/a
tank 0.454 0.766 0.427 0.429 0.891 n/a
bridge 0.441 0.885 0.368 0.411 0.609 0.714
harbor 0.711 0.979 0.858 0.601 0.751 0.756
Average 0.564 0.882 0.616 0.511 0.733 0.624